### v1

# Training Noise Token Pruning

Abstract
In the present work we present Training Noise Token (TNT) Pruning for vision transformers. Our method relaxes the discrete token dropping condition to continuous additive noise, providing smooth optimization in training, while retaining discrete dropping computational gains in deployment settings. We provide theoretical connections to Rate-Distortion literature, and empirical evaluations on the ImageNet dataset using ViT and DeiT architectures demonstrating TNT’s advantages over previous pruning methods.
Figure 1: Training Noise Token Pruning (TNT). Our proposed method computes a relevance term $\alpha_{i}$ for each token. In training (diagrammed at top), these terms dictate an amount of noise added to the token, while at test time they indicate pruning order.
1 Introduction
Token pruning is a class of methods for reducing computational load in transformers [20] by reducing the input length. While transformers are already somewhat robust to dropping a small number of tokens at random, learned dropping schemes enable larger dropping rates and thereby higher speed-ups with smaller accuracy penalties. These gains are only amplified by operations super-linear cost scaling (e.g., attention) and memory footprint.
Token pruning methods exploit predictive information differences between tokens; for a given task, some tokens are more useful than others. By dropping the least informative tokens first, a learned method can preserve the most amount of accuracy while removing desired number of tokens. While these methods were originally explored in a natural language processing context, token redundancy is arguably stronger in image transformers, and multiple token pruning methods for vision models (e.g. ViT [8]) have been proposed [17, 23, 21, 22, 15]. Unfortunately, token pruning as an action is discrete and not easily optimized directly. Current techniques use either forms of stochastic discrete dropping [17, 23], or heuristics [21] using attention activations themselves.
In this paper, we reframe Token Pruning as a constrained case of the Information Bottleneck problem. By viewing the token dropping rate as a channel constraint and the accuracy penalty as a distortion metric, we can directly apply ideas from the compression literature. Relaxations of these conditions provide continuous optimization desiderata, which are more easily optimized. Even though the original problem is difficult to solve exactly due to high-order token interaction information (“synergy” and “redundancy”), we show that a simple approximation to this case works surprisingly well. Specifically, we demonstrate that the proposed method has superior performance in comparison to recent pruning methods using either stochastic discrete dropping or attention heuristics on standard base models for the ImageNet-1K benchmark, in particular in the low token regime (i.e., when removing a large percentage of tokens).
In the present work we provide:
  • A novel method for token-pruning that provides state-of-the-art performance with respect to the accuracy/computation load trade-off.
  • A justification and intuition for that method based on the information bottleneck.
  • Empirical experiments demonstrating the use and utility of our method along with previous methods as baselines, with evaluations on ImageNet [8] using two common image transformers, ViT [8] and DeIT [19], as base architectures.
Our code can be found at https://github.com/mx-ethan-rao/tnt.
2 Related Work
Token Pruning:
Token pruning (or “token dropout”) has been explored for both transformer-based language models [10, 12, 13] and vision transformers [17, 22, 23, 21]. Broadly speaking, these methods can be separated into two categories: stochastic dropout methods, where tokens are randomly removed based on a per-token computed likelihood [10, 12, 13, 17, 23], and heuristic attention-based methods [15, 21]
Rao et al. 2021 [17] introduces Dynamic ViT, which is notable as the first token pruning method for vision transformer [8]. It is prototypical of the stochastic dropout family of methods: it defines a relaxation of the rate criterion and samples tokens according to an inclusion likelihood. Yin et al. 2022 [23] introduce a refinement of this method (Adaptive ViT), incorporating a halting module and associated score, as well as a loss function that encourages early stopping.
Along a different path, multiple heuristics have been introduced for token pruning based on the attention scores [15, 21]. These use the intuition that tokens to which other tokens ascribe high attention (highly attended tokens) are of high importance. Liang et al. 2022 [15] additionally merges the lowest attended tokens, while Wang et al. 2024 puts this both into a graph ranking framework (PageRank [16]) and into the zero-shot context.
A simple baseline version of an attention-based heuristic, “Top-K pruning”, was also found to perform competitively as well [11]. This method simply ranks tokens based on the attention distribution to a CLS token, and then truncates after the top $K$. While it is not directly applicable to architectures without CLS tokens or non-classification tasks, our results in Section 4 show that it outperforms most other methods where it can be applied.
Merger Methods:
Complementary to dropping methods, token mergers and similarity-based pruning methods also decrease transformer computational cost by reducing the size of the token set [5, 21, 15]. Similarity-based pruning [21] exploits exactly the opposite problem in the token set; instead of removing low relevance tokens they generally merge redundant tokens. Merger methods are more general, also possibly allowing for encoding of background context variables [5]. The use of these methods is not mutually exclusive with token pruning [21]. While in the present work we do not provide a completely novel merger method, we provide improvements on an existing similarity-based pruning step, and provide justification for its necessity.
Information Bottleneck:
Our framework for token dropping relies upon theory originally explored in the Information Bottleneck [18, 2], which in turn is based upon Rate-Distortion theory [3]. The information bottleneck characterizes encodings in terms of their relevance (the additive inverse of a distortion metric) and a rate constraint. We place token pruning’s two relevant metrics into this context, which then provides a natural relaxation for the rate constraint.
Figure 2: Noise Allocator block architecture: the block diagrammed above is injected into pre-trained models as a pruning layer. It takes the output of the previous Transformer block as input, then computes the noise signal terms $\alpha$ using a linear layer followed by a Softmax function. During training it samples Gaussian noise conditioned on $\alpha$ for each token, then adds the noise to the token embeddings. At test time, tokens are instead dropped. This pruning method can be trained with all parameters outside the noise allocator are frozen.
3 Methodology
3.1 Token pruning as bottleneck optimization
We model token pruning optimization as a set selection problem. Given a transformer architecture $f$, we want to select the token subset $T\subset\mathcal{T}$ that minimizes loss $\mathcal{L}$ for each datapoint and label pair $(x,y)$ subject to constraint $|T|<K$:
$\min_{T\subset\mathcal{T}}\mathcal{L}(f(x|_{T}),y)\text{ s.t. }|T|<K$ (1)
where $\mathcal{T}$ is the set of all tokens, and $x|_{T}$ are input data restricted to the token set $T$.
We can map this problem into the information-theoretic literature by drawing analogies between the loss and constraint functions in Eq. 1 and the rate and distortion functions of compression literature [18]. The loss function usually measures discrepancies between $y$ and its estimator $f(x|_{T})$, and, as in previous work [18, 2], is a direct analog to the distortion function.
The analog to the rate function requires more assumptions. While viewing each token as a fixed number of bits directly maps the $|T|<K$ constraint into a transmission rate, this does nothing to aid our optimization. Taking the mapping one step further, we could instead use the mutual information $I(x|_{T},x)<K$ constraint. This forces the effective information transmission to remain below $K$ bits. We then would like to solve
$\min_{T\subset\mathcal{T}}\text{Distortion}(f(x|_{T}),y)\text{ s.t. }I(x|_{T},x)<K$ (2)
which is exactly the Information Bottleneck.
This formulation is at once quite useful in that it maps a discrete optimization into a continuous problem, yet unfortunately also problematic in its exact computation. The rate parameter is definitionally aware of the content of the tokens. This provides us with a route to a potential procedure: force the token content to have at most $K$ bits of information from the original $x$ by adding an external source of noise. Selecting what information to keep from $x$ is equivalent to token selection, but still enjoys a continuous optimization landscape, so long as we choose to add continuous noise. On the otherhand, in order to exactly measure the mutual information we need to measure redundancy between tokens. This issue also was implicitly present in the discrete form of the optimization problem, but it is explicitly included in mutual information.
3.2 Proposed approximate solution
Solutions to the information bottleneck have been proposed in the literature, both in a general setting [18] and in a deep learning context [1, 2]. However, for token pruning, these solutions are almost certainly too computationally heavy to be efficient as dropping methods. Instead, we will solve a simplified version that ignores redundant information, then prune redundant tokens in a second phase. Surprisingly, even with the errors induced by approximation, this method produces accurate results with very little compute overhead.
Given a non-pruned vision transformer architecture, we train a module we call the Noise Allocator to estimate relative token relevance scalars $\alpha_{i}$ for each token. Due to the residual connections used by many transformer architectures, this module can be placed after any of the individual transformer blocks, or after multiple blocks (in a multi-layer dropping case).
Consider an architecture with L repeated transformer blocks possibly after an initial token embedding layer (for example, ViT has 12 identical blocks). Let the $\ell$th Transformer block $f^{(\ell)}$ with input $x\in\mathbb{R}^{N\times D}$, where $N$ is the number of tokens, and $D$ is the embedding dimension. Let $x^{(\ell)}_{i}$ denote the $i$th token. For every block where the user would like to perform pruning, we introduce a linear token relevance predictor with a weight matrix $W^{(\ell)}\in\mathbb{R}^{D\times 1}$.
We then compute $\alpha_{i}$ as
$\alpha_{i}^{\ell}=\text{Softmax}(W^{(\ell)}f^{\ell}(x^{(\ell)}))$ (3)
For training/fine-tuning, we compute noise variables $\eta_{i}^{\ell}$ as
$\eta_{i}^{\ell}=(1-\alpha_{i}^{(\ell)})\varepsilon,$ (4)
where $\varepsilon\sim\mathcal{N}(0,\beta I_{D})$ (“the reparameterization trick” [14]). Here, $\beta$ is a hyper-parameter that controls the total amplitude of noise added, and corresponds with $K$ up to a monotonic transformation. The Softmax function in Eq. 3 ensures that the network must allocate $\beta$ amount of noise; without it, a degenerate zero-noise solution exists.
Classical results in Information Theory state that the mutual information (channel capacity) for each token to its noised variant has an upper bound proportional to $\log(1+P_{signal}/P_{noise})$, where $P_{signal}/P_{noise}$ is the ratio of the power (amplitude) of the token signal and to that of the noise [6] (the signal to noise ratio). Due to the layer-wise normalization (LayerNorm) of many transformer architectures, $P_{signal}$ is necessarily bounded. Disregarding the synergistic information between tokens, optimizing our proposed Noise Allocator amounts to solving an information bottleneck problem. We discuss limitations of this approach in Section 4.4.
At test time noise is not sampled, and instead of $\beta$ we directly select the top $K$ tokens, dropping the rest of the token set. We interpret $\alpha^{(\ell)}$ as a token importance map for $x^{(\ell)}$. Based on this importance ranking, we may either discard less important tokens or merge them with other tokens in the subsequent stage.
Weight vector $W$ is applied to every $x_{i}$ identically, so the additional parameter space complexity remains $\mathcal{O}(NL)$, where $L$ is the number of layers at which tokens are selectively dropped during testing. Given $L\ll D$, this complexity effectively reduces to $\mathcal{O}(D)$, making it comparatively negligible within the overall model size. For weaker encoders and/or smaller transformer blocks such as DeiT-Tiny, it is sometimes helpful to increase the complexity of $W$ to a multi-layer structure (e.g., an MLP), as we show in experiments in the Appendix.
3.2.1 Similarity-based Pruning by Random Partition (Redundancy Removal)
Even though individual tokens might be highly relevant to the classification problem (i.e, individually $I(x_{i}^{\ell},y)$ might be high), relevant tokens sharing a high amount of information with a kept token represents wasted capacity. In order to avoid this waste, we implement a redundancy removal step for test time usage, which removes similar and therefore possibly redundant tokens.
The method is as follows: Tokens are randomly divided into two groups. We then identify the closest matching token in one group for each token in the other, recording the similarity scores of each paired token based on their token embeddings. Next, we prune the top-$r$ most similar pairs and prune the associated tokens in the second group.
Our similarity-based pruning approach closely resembles that of Zero-TP [21], with two key modifications. First, rather than using the Key values as the partitioning metric, we directly use $x^{(l)}$ so that it can be applied directly after the Noise Allocator step.Second, instead of sequentially partitioning tokens based on their importance scores, we apply a random partitioning strategy. Our ablation study suggests that random partitioning yields improved performance for the proposed model.
Figure 3: Visualization of Token Pruning maps on ImageNet-1K: at left are the original images, and at each column progressing right are single layer prunings and their associated kept/dropped tokens, for layers 1-5 of the DeiT-B-Distil. model.
Figure 4: Single Layer Pruning results: We plot the Top-1 Accuracy in the ImageNet-1k validation set for each of the pruning methods as a function of computational efficiency, in the top row measured by GFLOPs and in the bottom row measured by throughput, for single layer pruning. The base model is DeiT-B-Distil in the first column, DeiT-S-Distil. in the second column, and ViT/16 in the third column.
K=1.0 K=0.8 K=0.6 K=0.5 K=0.25
#tokens=196 #tokens=156 #tokens=127 #tokens=98 #tokens=49
DeiT-S-Distil. Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs
Random Drop 80.50 4.63 79.54 3.90 77.83 3.20 76.57 2.87 67.77 2.03
ToMe [5] 80.07 3.85 77.40 3.11 73.13 2.75 - -
Zero-TP [21] 80.00 3.91 78.94 3.21 77.74 2.88 69.99 2.05
DynamicViT [17] 79.39 3.99 78.66 3.28 77.96 2.94 71.03 2.08
Top-K [22] 80.12 3.85 78.89 3.11 77.69 2.75 68.57 1.86
TNT (ours) 80.11 3.90 79.29 3.20 78.65 2.87 72.66 2.03
#tokens=196 #tokens=160 #tokens=130 #tokens=100 #tokens=60
ViT/16 Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs
Random Drop 78.70 9.17 77.17 7.69 75.12 6.50 69.21 5.33 31.16 3.81
ToMe [5] 77.71 7.66 75.70 6.42 66.46 5.22 - -
Zero-TP [21] 77.43 7.32 75.96 6.52 70.92 5.35 40.84 3.83
TNT (ours) 77.94 7.70 76.22 6.50 71.69 5.33 41.75 3.81
Table 1: Top-1 Accuracy on the ImageNet-1K validation set (Acc.) and computational cost (GFLOPs) across methods for differing keep rates $K$, using a single layer of token pruning. The left most column are the performance and computational cost of the base architectures.
Figure 5: Multi-layer Pruning results: We plot the Top-1 Accuracy in the ImageNet-1k validation set for each of the pruning methods as a function of computational efficiency, in the top row measured by GFLOPs and in the bottom row measured by throughput, for multi-layer pruning. The base model is DeiT-B-Distil in the first column, DeiT-S-Distil. in the second column, and ViT/16 in the third column.
Method Acc. GFLOPs TP (imgs/s)
High Token Regime GFLOPs $\geq 8.0$ Deit-B-Distil. [19] 82.55 17.68 379
Top-K [15] 81.82 13.18 504
Zero-TP [21] 81.92 13.10 433
DynamicViT [17] 80.91 13.34 498
ToMe [5] 81.86 12.11 508
TNT (ours) 81.97 13.11 509
GFLOPs $\geq 2.0$ Deit-S-Distil. [19] 80.49 4.63 1150
Top-K [15] 79.81 3.44 1496
Zero-TP [21] 79.66 3.42 983
DynamicViT [17] 79.45 3.47 1448
ToMe [5] 80.12 3.45 1281
TNT (ours) 79.89 3.41 1443
GFLOPs $\geq 4.0$ ViT/16 [8] 78.70 9.17 644
Zero-TP [21] 77.07 6.75 688
ToMe [5] 77.03 6.97 734
TNT (ours) 77.32 6.73 842
Low Token Regime GFLOPs $<8.0$ Deit-B-Distil. [19] 82.55 17.68 379
Top-K [15] 55.98 5.93 1084
Zero-TP [21] 19.76 6.18 731
DynamicViT [17] 11.20 5.79 1102
ToMe [5] 43.37 5.89 995
TNT (ours) 59.93 5.87 1095
GFLOPs $<2.0$ Deit-S-Distil. [19] 80.49 4.63 1150
Top-K [15] 58.20 1.57 3003
Zero-TP [21] 27.75 1.63 1485
DynamicViT [17] 24.51 1.52 3037
ToMe [5] 62.88 1.55 2486
TNT (ours) 63.82 1.53 2973
GFLOPs $<4.0$ ViT/16 [8] 78.70 9.17 644
Zero-TP [21] 0.63 3.26 1179
ToMe [5] 6.30 3.26 1485
TNT (ours) 13.42 2.98 1786
Table 2: Top-1 Accuracy on the ImageNet-1K validation set (Acc.) and computation cost measured by GFLOPs and throughput (TP), for DeiT-B-Distil., DeiT-S-Distil., and ViT/16, in two different computing regimes (“High” and “Low”, defined for each base model).
4 Experiments
We conduct a series of experiments on pre-trained vision transformer models, including DeiT (Tiny [^1] DeiT-Tiny results relegated to the Appendix., Small, Base) [19] and ViT/16 [8], each trained on ImageNet-1K [7]. We then compare our model with previously proposed token-dropping models, as well as two simple baseline schemes: DynamicViT [17], ZeroTP[21], ToMe [5], Top-K [11], and random dropping. In Section 4.3, we present ablation studies to validate our individual design choices.
We first show qualitative results for pruning $50\%$ of tokens at varying layers (single-layer pruning at layers 1-5) on the ImageNet-1K validation dataset in Figure  3. As seen at left in the figure, extremely early layers result in relevant tokens being dropped, and irrelevant tokens being included; this matches with results from previous literature [17, 9, 21]. Results further indicate that the the importance score based on noise signal term $\alpha$ becomes more reasonable in latter layers, performing well on a variety of qualitatively different images (tiny objects, mixed foreground and background, low contrast separation). Additional examples are provided in the Appendix.
We conduct a comprehensive evaluation for each method by sweeping the token keep rate (K) from low to high-token regime, as detailed in Sections 4.1 and 4.2. We experiment with both single-layer and multi-layer pruning schema. In the single-layer scheme all token pruning occurs at a single layer, usually early in the network. In comparison, the multi-layer scheme has pruning blocks spread across the architecture. The multi-layer scheme is generally more effective, albeit with additional computational cost, but due to the large number of possible parameters for multi-layer pruning introduced by the multiple keep rates (with possibly varying predictive accuracy), may be hard to effectively tune for all methods. This makes comparisons in the multi-layer scheme inexact, which is why we include the single layer experiments. We evaluate each model in terms of Top-1 Accuracy, FLOPs, and throughput. Our model achieves top performance across most experiments and remains competitive on others.
Experimental Setup: All training and evaluations are conducted on a single compute node equipped with 8 NVIDIA A40 GPUs. The experiments are performed on the ImageNet-1K dataset [7], with all images resized to a resolution of 224px. We evaluate the proposed method using only frozen pre-trained backbone methods, training the Noise Allocator weights for 40 epochs. And we set $\beta$ (a hyperparameter controlling the magnitude of added noise) to 0.02 for all base models while training. For testing, a single GPU to measure the computational performance.
4.1 Rate Sweep for Single Layer Pruning
The experiments of previous studies have shown that most models [17, 9, 21] refrain from dropping tokens in initial Transformer layers, as these layers contain limited information relevant to token importance. For our experiments, we select the second layer for ViT and third layer for DeiT variants as the standard layers for single-layer token pruning. We use the standard pre-trained DeiT variants for all experiments for all models. For all ViT experiments for all models, we use a modified ViT/16 with smaller embedding dimension $mlp\_dim=emb\_dim=768$. Additionally, we use mean-pooled token embeddings for predictions instead of the CLS token.
Plots for pruning performance on DeiT-S-Dist., DeiT-B-Dist., ViT/16 across different pruning models are presented in Figure 4, for both Top-1 Accuracy versus GFLOPs and Top-1 Accuracy versus throughput (images per second). In general throughput is not one-to-one with GFLOPs due to differing parallelism between models. Performance between all methods converge to the base model case as the keep rate approaches 1.0. For decreasing $K$ however, the proposed model shows superior performance until $K\leq 0.25$, at which point baseline models overtake the proposed model, albeit with all models experiencing strong performance degradation. In general our results for the baseline models mirror that of Haurum et al. 2023 [11], and similar to their report we also find that the Top-K baseline is generally the strongest where it can be applied (i.e., for models incorporating the CLS token), aside from our proposed method.
Overall, the number of tokens pruned is consistent across models for a given keep rate. For TNT, we select top $(NK+s)$ at the Noise Allocator stage and prune $s$ tokens at the similarity pruning stage for a given keep rate $K$ and total number of tokens $N$. We set $s$ to be 25 and 30 for experiments involving DeiT and ViT, respectively. We apply the same setting to Zero-TP [21], as it also utilizes similarity-based pruning. We could not reproduce the results of Yin et al. [23] in our own experimental context. In Table 1, we report specific numerical results for ViT/16 and DeiT-S-Distl, contingent on keep rates $K$, instead of GFLOPs or Throughput as in the plots. Complete numerical results and details such as keep rates for each experiment are also provided in the Appendix. For ToMe [5], it is impossible to prune more than 50% of tokens within a single layer, so lower single-layer pruning is not possible.
4.2 Rate Sweep at Multi-layer Pruning
We evaluate the performance of token pruning across multiple layers as this better reflects a performant pruning system, with the caveat that equal comparison conditions are more difficult to ensure. For the token keep rate at each layer and the specific layers chosen for token pruning (i.e. pruning locations), we strictly follow the instructions on the original paper for each model. For Zero-TP [21], tokens are pruned at layers [1, 3, 6, 9, 11], where layers [1, 11] perform similarity-based pruning only; for ToMe [5], pruning occurs at every layer; and for DynamicViT [17], pruning is applied at layers [3, 6, 9]. As Top-K lacks specific pruning instructions, we align its pruning layers with those of TNT. We gradually sweep the keep rate at pruning locations to generate Figure 5 for DeiT-B-Distil., DeiT-S-Distil., and ViT/16. We also provide partial numerical results in Table 2. All base models are the same as those used in the single-layer pruning section. Additional details on experimental setups and specific details for each model are provided in the Appendix.
For TNT, we prefer pruning tokens at earlier layers, as it provides stronger performance, and as shown in Figure 3, layers [3, 4, 5] serve as effective locations for token pruning. Rather than performing the similarity pruning stage at every pruning layer as in Zero-TP, we prune $s$ tokens before tokens are sent to the ViT blocks. Our ablation study justify this. For all experiments, $s$ is set to 40. Our results indicate that TNT consistently shows strong performance in varying token (computation) ranges. As the proportion of pruned tokens increases, the importance of the retained tokens becomes more critical. TNT shows a larger performance gap compared to other models in the low-token regime.
4.3 Ablation Study
We conducted an ablation experiment to justify inclusion of the similarity-based pruning by random partition, as opposed to token merger or sequential partition. We find a slight boost to performance with similarity based pruning, with a very minor bias towards the random partition method. Previous works [5, 4] explore token merging, which can be viewed as another form of pruning. For our method, this approach did not give an improvement over similarity pruning.
Method Acc. GFLOPs TP (imgs/s)
Deit-B-Distil. [19](SL) 82.55 17.68 379
TNT w/o Sim. Pruning 81.33 12.29 546
TNT (Seq. Part.) 81.52 12.30 536
TNT (Token Merge) 81.54 12.30 538
TNT (Random Part.) 81.57 12.30 542
Deit-B-Distil. [19](ML) 82.55 17.68 379
TNT w/o Sim. Pruning 80.55 11.03 605
TNT (Token Merge) 81.41 11.41 575
TNT (Random Part.) 81.50 11.41 581
Table 3: Ablation study for TNT in DeiT-S-Distil. “SL” is single-layer pruning; “ML” is multi-layer pruning; Acc. is Top-1 Accuracy, TP is the Throughput, measured in images-per-second.
4.4 Limitations
While the proposed TNT method provides relatively improved performance over other methods, important gaps remain in methodology. In this method, redundant tokens are not removed during training; while this can be performed through merging as in ToMe [5], as experimental results show it is difficult to do in a stable manner. Further, synergistic information between tokens is not considered; this seems less relevant to ImageNet classification, but likely would be useful in more complicated label structures (e.g., hierarchical/multi-class structures). This is also perhaps a deeper problem than redundant tokens, as it has combinatorial complexity in the order of the interactions considered.
Beyond this, for better comparison measurements multi-layer dropping should be tuned for each method. This is prohibitively expensive, but could provide some less stable but higher capacity methods with a boost in performance.
Finally, in a deployment setting, hardware constraints will dictate keep-rates, which could be optimized for in the base-models directly (i.e., we could optimize a ViT for a $50\%$ keep rate, or for a specific memory structure). This optimization was not done here, nor do we propose any candidate deployment device constraints, but nevertheless should be considered for best performance.
5 Conclusion
In this work, we introduced a novel token pruning method within the Information Bottleneck framework, aiming to optimize vision transformers by continuous optimization of token pruning. Our extensive evaluations on the ImageNet dataset using ViT and DeiT architectures demonstrate state-of-the-art performance in the accuracy-computation trade-off compared to existing methods. Specifically, our method excels in low-token retention rates, maintaining high accuracy while significantly reducing computational loads. These results underscore the potential impact of our method in improving the efficiency of deploying vision transformers, particularly in resource-constrained applications.
Supplementary Material
A Implementation Code
Figure 6 is an implementation of our “VisionTransformerWithTNT” in PyTorch.
1class VisionTransformerWithTNT(VisionTransformer):
2 def __init__(self, *args, **kwargs):
3 super().__init__(*args, **kwargs)
4 # Parameters introduced: Add alpha heads to produce noise signal term
5 self.alpha_norm = kwargs[’norm_layer’](self.embed_dim)
6 self.alpha_heads = nn.ModuleList([
7 nn.Linear(self.embed_dim, 1) for _ in range(kwargs[’depth’])
8 ])
9 self.alpha_heads.apply(self._init_weights)
10
11 def forward_features(self, x):
12 B = x.shape[0]
13 x = self.patch_embed(x)
14
15 cls_tokens = self.cls_token.expand(B, -1, -1)
16 x = torch.cat((cls_tokens, x), dim=1)
17 x = x + self.pos_embed
18 x = self.pos_drop(x)
19 for i, (blk, alpha_head) in enumerate(zip(self.blocks, self.alpha_heads)):
20 x = blk(x)
21 # Noise allocator: To add noise to token embeddings at 1-5 layers while fine-tuning
22 if self.training and i < 5:
23 alpha = alpha_head(x[:, 1:])
24 alpha = torch.softmax(alpha.squeeze(-1), dim=-1)
25 alpha = 1 - alpha
26 noise = torch.randn_like(x[:, 1:]) * alpha.unsqueeze(-1).repeat(1,1,x.size(-1))
27 zero_noise = torch.zeros_like(cls_tokens)
28 noise = torch.cat((zero_noise, noise), dim=1)
29 x = self.alpha_norm(x)
30 x = x + 0.02 * noise
31 x = self.norm(x)
32 return x[:, 0]
33
34 def forward(self, x):
35 x = self.forward_features(x)
36 x = self.head(x)
37 return x
Figure 6: Python implementation of VisionTransformerWithTNT class. Codes highlighted with brown are the main modifications. VisionTransformer class is taken from https://github.com/rwightman/pytorch-image-models/blob/master/timm/models/vision_transformer.py. We make simple modifications to allocate noise to token embeddings while fine-tuning.
B Additional Experiment Details
Full results for plots and tables in the main paper including DeiT-Tiny-Distil. [19]. For all results, Acc. is Top-1 Accuracy, TP is the Throughput, measured in images-per-second.
Single-layer pruning settings: K is the keep rate. We divide K values into two tables: K = 0.8 to 0.45 represent the high-token regime, while K = 0.4 to 0.2 correspond to the low-token regime. Highlighted rows achieve the highest accuracy in a specific token regime. For Top-K [15], GFLOPs is relatively smaller as pruning occurs in the middle of the Transformer block. For clarity in visualizations of the computation-accuracy trade-off, data of Top-K corresponding to K=0.2 is omitted from the plots but is reported in the table.
Multi-layer pruning settings: For TNT, Top-K [15], and Zero-TP [21], Loc and Rate denote the pruning locations (layers) and the pruning rate at each layer, respectively. itr is an additional parameter for Zero-TP only, which specifies the number of iterations for the PageRank algorithm at each pruning layer. For ToMe [5], pruning is performed at every layer, with $\bm{r}$ representing the number of tokens pruned per layer. To ensure GFLOPs alignment for better comparison, we only collect 8 data points per base model for ToMe (10 for other models). For DynamicViT [17], the keep rates for the three pruning layers are specified as [$\bm{\rho}$, $\bm{\rho}^{2}$, $\bm{\rho}^{3}$]. Rate, $\bm{r}$, and $\bm{\rho}$ vary across different experiments.
Table 4 provides an overview of the fixed hyperparameters used consistently in all experiments. For different base models, we conduct the hyperexperiments using the same set of parameters.
Method Param.
Top-K [15] loc=[3, 4, 5]
Zero-TP [21] itr(single-layer)=50
loc=[1, 3, 6, 9, 11]
itr(multi-layer)=[30, 5, 5, 1, 1]
DynamicViT [17] loc=[3, 6, 9]
ToMe [5] loc[1:12]
TNT (ours) loc=[3, 4, 5]
Table 4: overview of the fixed hyperparameters used consistently in all experiments.
B.1 Base Model: DeiT-Base-Distil.
For single-layer pruning, high-token regime results are listed in Table 5; low-token regime results are listed in Table 6. For multi-layer pruning, high-token regime results are listed in Table 7; low-token regime results are listed in Table 8.
Method Acc. GFLOPs TP (imgs/s)
Deit-B-Distil. 82.55 17.68 379
K = 0.8 Random Drop 81.59 14.93 -
Top-K [15] 82.16 14.74 462
Zero-TP [21] 82.25 14.95 411
DynamicViT [17] 80.87 15.33 438
ToMe [5] 81.86 14.74 447
TNT (ours) 82.31 14.93 455
K = 0.7 Random Drop 80.79 13.64 -
Top-K [15] 81.63 13.36 507
Zero-TP [21] 81.74 13.67 448
DynamicViT [17] 80.66 14.00 481
ToMe [5] 80.80 13.36 493
TNT (ours) 82.05 13.64 495
K = 0.6 Random Drop 79.59 12.29 -
Top-K [15] 80.70 11.92 558
Zero-TP [21] 80.58 12.32 488
DynamicViT [17] 80.02 12.61 526
ToMe [5] 78.28 11.92 546
TNT (ours) 81.57 12.30 542
K = 0.5 Random Drop 77.72 11.02 -
Top-K [15] 79.25 10.56 635
Zero-TP [21] 78.63 11.05 540
DynamicViT [17] 79.06 11.31 588
ToMe [5] 71.57 10.56 621
TNT (ours) 80.62 11.03 605
K = 0.45 Random Drop 76.23 10.35 -
Top-K [15] 78.06 9.85 677
Zero-TP [21] 77.15 10.38 571
DynamicViT [17] 78.26 10.63 624
TNT (ours) 79.85 10.31 640
Table 5: Single-layer pruning for DeiT-B-Distil. in high-token regime.
Method Acc. GFLOPs TP (imgs/s)
Deit-B-Distil. 82.55 17.68 379
K = 0.4 Random Drop 74.08 9.70 -
Top-K [15] 76.53 9.14 728
Zero-TP [21] 75.15 9.72 603
DynamicViT [17] 77.01 9.95 668
TNT (ours) 78.89 9.70 684
K = 0.35 Random Drop 70.57 9.04 -
Top-K [15] 74.26 8.43 786
Zero-TP [21] 72.32 9.06 649
DynamicViT [17] 75.28 9.27 710
TNT (ours) 77.22 9.04 723
K = 0.3 Random Drop 64.89 8.38 -
Top-K [15] 70.68 7.73 853
Zero-TP [21] 68.68 8.41 680
DynamicViT [17] 72.69 8.59 770
TNT (ours) 74.75 8.38 786
K = 0.25 Random Drop 57.12 7.80 -
Top-K [15] 65.95 7.10 925
Zero-TP [21] 63.62 7.82 717
DynamicViT [17] 68.87 7.99 823
TNT (ours) 70.76 7.80 834
K = 0.2 Random Drop 44.17 7.14 -
Top-K [15] 57.51 6.40 1016
Zero-TP [21] 55.28 7.17 779
DynamicViT [17] 61.86 7.32 904
TNT (ours) 62.62 7.15 907
Table 6: Single-layer pruning for DeiT-B-Distil. in low-token regime.
Method Acc. GFLOPs TP (imgs/s) Param.
Deit-B-Distil. 82.55 17.68 379 -
GFLOPs $\approx 13.0$ Top-K 81.82 13.18 503 Rate=[.9, .9, .8]
Zero-TP [21] 81.92 13.10 433 Rate=[1., .9, .9, .9, 1.]
DynamicViT [17] 80.91 13.34 498 $\rho$=.8
ToMe [5] 81.86 12.11 508 $r$=8
TNT (ours) 81.97 13.11 509 Rate=[1., .95, .95]
GFLOPs $\approx 11.5$ Top-K 80.96 11.72 581 Rate=[.85, .8, .8]
Zero-TP [21] 80.97 11.37 424 Rate=[1., .8, .8, .9, 1.]
DynamicViT [17] 80.67 11.49 576 $\rho$=.7
ToMe [5] 81.71 11.56 533 $r$=11
TNT (ours) 81.50 11.41 581 Rate=[.9, .9, .9]
GFLOPs $\approx 10.0$ Top-K 79.63 10.25 650 Rate=[.85, .7, .7]
Zero-TP [21] 78.77 9.71 573 Rate=[1., .7, .7, .8, 1.]
DynamicViT [17] 79.55 9.88 659 $\rho$=.6
ToMe [5] 80.51 9.39 651 $r$=15
TNT (ours) 80.56 10.01 660 Rate=[.85, .85, .8]
GFLOPs $\approx 8.5$ Top-K 76.84 8.75 748 Rate=[.7, .7, .65]
Zero-TP [21] 75.60 8.61 613 Rate=[1., .6, .7, .7, 1.]
DynamicViT [17] 76.73 8.57 761 $\rho$=.5
TNT (ours) 78.96 8.77 758 Rate=[.75, .85, .7]
GFLOPs $\approx 8.0$ Top-K 75.04 8.14 814 Rate=[.65, .65, .65]
Zero-TP [21] 73.32 8.22 636 Rate=[1., .6, .6, .7, 1.]
DynamicViT [17] 73.71 7.96 817 $\rho$=.45
TNT (ours) 77.31 8.03 822 Rate=[.65, .85, .7]
Table 7: Multi-layer pruning for DeiT-B-Distil. in high-token regime.
Method Acc. GFLOPs TP (imgs/s) Param.
Deit-B-Distil. 82.55 17.68 379 -
GFLOPs $\approx 7.5$ Top-K 71.59 7.37 881 Rate=[.6, .6, .6]
Zero-TP [21] 69.66 7.76 654 Rate=[1., .5, .5, .6, 1.]
DynamicViT [17] 68.79 7.45 877 $\rho$=.4
ToMe [5] 72.29 7.23 823 $r$=20
TNT (ours) 75.33 7.49 871 Rate=[.6, .8, .7]
GFLOPs $\approx 7.0$ Top-K 67.51 6.84 948 Rate=[.6, .5, .6]
Zero-TP [21] 55.65 7.03 703 Rate=[1., .4, .4, .4, 1.]
DynamicViT [17] 60.30 6.97 926 $\rho$=.35
ToMe [5] 68.53 6.91 855 $r$=21
TNT (ours) 71.92 6.89 947 Rate=[.5, .8, .7]
GFLOPs $\approx 6.6$ Top-K 65.24 6.56 993 Rate=[.55, .5, .6]
Zero-TP [21] 43.24 6.79 735 Rate=[1., .4, .3, .4, 1.]
DynamicViT [17] 45.65 6.53 982 $\rho$=.3
ToMe [5] 64.33 6.62 893 $r$=22
TNT (ours) 69.06 6.53 991 Rate=[.5, .7, .7]
GFLOPs $\approx 6.2$ Top-K 61.54 6.29 1044 Rate=[.55, .5, .5]
Zero-TP [21] 39.17 6.37 762 Rate=[1., .3, .4, .4, 1.]
DynamicViT [17] 29.87 6.17 1043 $\rho$=.25
ToMe [5] 51.32 6.11 962 $r$=24
TNT (ours) 66.21 6.28 1039 Rate=[.5, .7, .6]
GFLOPs $\approx 5.7$ Top-K 55.98 5.93 1084 Rate=[.5, .45, .5]
Zero-TP [21] 19.76 6.18 731 Rate=[1., .3, .3, .4, 1.]
DynamicViT [17] 11.20 5.79 1102 $\rho$=.2
ToMe [5] 43.37 5.89 995 $r$=25
TNT (ours) 59.93 5.87 1095 Rate=[.5, .6, .55]
Table 8: Multi-layer pruning for DeiT-B-Distil. in low-token regime.
B.2 Base Model: DeiT-Small-Distil.
For single-layer pruning, high-token regime results are listed in Table 9; low-token regime results are listed in Table 10. For multi-layer pruning, high-token regime results are listed in Table 11; low-token regime results are listed in Table 12.
Method Acc. GFLOPs TP (imgs/s)
Deit-S-Distil. 80.49 4.63 1150
K = 0.8 Random Drop 79.54 3.90 -
Top-K [15] 80.12 3.85 1338
Zero-TP [21] 79.99 3.91 1156
DynamicViT [17] 79.39 3.99 1255
ToMe [5] 80.07 3.85 1290
TNT (ours) 80.11 3.90 1316
K = 0.7 Random Drop 78.91 3.55 -
Top-K [15] 79.69 3.48 1478
Zero-TP [21] 79.55 3.57 1243
DynamicViT [17] 79.25 3.64 1381
ToMe [5] 79.24 3.49 1415
TNT (ours) 79.82 3.56 1427
K = 0.6 Random Drop 77.83 3.20 -
Top-K [15] 78.89 3.11 1655
Zero-TP [21] 78.94 3.21 1343
DynamicViT [17] 78.66 3.28 1528
ToMe [5] 77.40 3.11 1596
TNT (ours) 79.29 3.20 1587
K = 0.5 Random Drop 76.57 2.87 -
Top-K [15] 77.69 2.75 1854
Zero-TP [21] 77.74 2.88 1463
DynamicViT [17] 77.96 2.94 1706
ToMe [5] 73.13 2.75 1791
TNT (ours) 78.65 2.87 1768
K = 0.45 Random Drop 75.61 2.69 -
Top-K [15] 76.70 2.57 1995
Zero-TP [21] 76.97 2.71 1514
DynamicViT [17] 77.22 2.76 1800
TNT (ours) 78.06 2.70 1878
Table 9: Single-layer pruning for DeiT-S-Distil. in high-token regime.
Method Acc. GFLOPs TP (imgs/s)
Deit-S-Distil. 80.49 4.63 1150
K = 0.4 Random Drop 74.57 2.52 -
Top-K [15] 75.45 2.38 2143
Zero-TP [21] 75.91 2.54 1605
DynamicViT [17] 76.40 2.58 1934
TNT (ours) 77.32 2.52 1994
K = 0.35 Random Drop 72.79 2.35 -
Top-K [15] 73.87 2.20 2342
Zero-TP [21] 74.58 2.37 1679
DynamicViT [17] 75.22 2.41 2091
TNT (ours) 76.19 2.35 2167
K = 0.3 Random Drop 70.42 2.18 -
Top-K [15] 71.60 2.02 2465
Zero-TP [21] 72.70 2.20 1711
DynamicViT [17] 73.41 2.24 2189
TNT (ours) 74.65 2.19 2269
K = 0.25 Random Drop 67.77 2.03 -
Top-K [15] 68.57 1.86 2713
Zero-TP [21] 69.99 2.05 1757
DynamicViT [17] 71.03 2.08 2376
TNT (ours) 72.66 2.03 2444
K = 0.2 Random Drop 62.83 1.87 -
Top-K [15] 64.12 1.68 2842
Zero-TP [21] 65.90 1.88 1797
DynamicViT [17] 66.67 1.91 2496
TNT (ours) 69.08 1.87 2552
Table 10: Single-layer pruning for DeiT-S-Distil. in low-token regime.
Method Acc. GFLOPs TP (imgs/s) Param.
Deit-S-Distil. 80.49 4.63 1150 -
GFLOPs $\approx 3.45$ Top-K [15] 79.81 3.44 1496 Rate=[.9, .9, .8]
Zero-TP [21] 79.66 3.42 983 Rate=[1., .9, .9, .9, 1.]
DynamicViT [17] 79.45 3.47 1448 $\rho$=.8
ToMe [5] 80.12 3.45 1281 $r$=8
TNT (ours) 79.89 3.41 1444 Rate=[1., .95, .95]
GFLOPs $\approx 3.0$ Top-K [15] 79.11 3.06 1706 Rate=[.85, .8, .8]
Zero-TP [21] 78.75 2.97 1171 Rate=[1., .8, .8, .9, 1.]
DynamicViT [17] 79.18 2.99 1679 $\rho$=.7
ToMe [5] 79.86 3.02 1448 $r$=11
TNT (ours) 79.38 2.97 1652 Rate=[.9, .9, .9]
GFLOPs $\approx 2.6$ Top-K [15] 77.65 2.67 1909 Rate=[.85, .7, .7]
Zero-TP [21] 76.92 2.54 1352 Rate=[1., .7, .7, .8, 1.]
DynamicViT [17] 78.38 2.57 1929 $\rho$=.6
ToMe [5] 79.12 2.54 1754 $r$=15
TNT (ours) 78.57 2.60 1882 Rate=[.85, .85, .8]
GFLOPs $\approx 2.25$ Top-K [15] 75.33 2.29 2190 Rate=[.7, .7, .65]
Zero-TP [21] 74.35 2.25 1389 Rate=[1., .6, .7, .7, 1.]
DynamicViT [17] 76.39 2.23 2192 $\rho$=.5
TNT (ours) 77.25 2.28 2143 Rate=[.75, .85, .7]
GFLOPs $\approx 2.1$ Top-K [15] 73.73 2.13 2343 Rate=[.65, .65, .65]
Zero-TP [21] 72.55 2.16 1415 Rate=[1., .6, .6, .7, 1.]
DynamicViT [17] 74.41 2.08 2321 $\rho$=.45
TNT (ours) 75.93 2.09 2302 Rate=[.65, .85, .7]
Table 11: Multi-layer pruning for DeiT-S-Distil. in high-token regime.
Method Acc. GFLOPs TP (imgs/s) Param.
Deit-S-Distil. 80.49 4.63 1150 -
GFLOPs $\approx 1.95$ Top-K 70.78 1.93 2525 Rate=[.6, .6, .6]
Zero-TP [21] 69.48 2.03 1357 Rate=[1., .5, .5, .6, 1.]
DynamicViT [17] 71.69 1.95 2490 $\rho$=.4
ToMe [5] 74.76 1.90 2176 $r$=20
TNT (ours) 74.54 1.95 2484 Rate=[.6, .8, .7]
GFLOPs $\approx 1.8$ Top-K 67.34 1.80 2701 Rate=[.6, .5, .6]
Zero-TP [21] 59.29 1.85 1331 Rate=[1., .4, .4, .4, 1.]
DynamicViT [17] 67.03 1.83 2635 $\rho$=.35
ToMe [5] 71.50 1.74 2351 $r$=21
TNT (ours) 72.26 1.79 2639 Rate=[.5, .8, .7]
GFLOPs $\approx 1.75$ Top-K 65.55 1.73 2794 Rate=[.55, .5, .6]
Zero-TP [21] 49.70 1.79 1471 Rate=[1., .4, .3, .4, 1.]
DynamicViT [17] 58.70 1.71 2761 $\rho$=.3
ToMe [5] 69.76 1.68 2377 $r$=22
TNT (ours) 70.31 1.70 2725 Rate=[.5, .7, .7]
GFLOPs $\approx 1.65$ Top-K 62.41 1.66 2896 Rate=[.55, .5, .5]
Zero-TP [21] 48.44 1.68 1481 Rate=[1., .3, .4, .4, 1.]
DynamicViT [17] 47.71 1.62 2943 $\rho$=.25
ToMe [5] 66.30 1.61 2357 $r$=24
TNT (ours) 68.36 1.64 2850 Rate=[.5, .7, .6]
GFLOPs $\approx 1.55$ Top-K 58.20 1.57 3003 Rate=[.5, .45, .5]
Zero-TP [21] 27.75 1.63 1485 Rate=[1., .3, .3, .4, 1.]
DynamicViT [17] 24.51 1.52 3037 $\rho$=.2
ToMe [5] 62.88 1.55 2486 $r$=25
TNT (ours) 63.82 1.53 2973 Rate=[.5, .6, .55]
Table 12: Multi-layer pruning for DeiT-S-Distil. in low-token regime.
B.3 Base Model: DeiT-Tiny-Distil.
Figure 7 shows the plots for DeiT-Tiny-Distil.. For single-layer pruning, high-token regime results are listed in Table 13; low-token regime results are listed in Table 14. For multi-layer pruning, high-token regime results are listed in Table 15; low-token regime results are listed in Table 16.
Figure 7: Experimental Results for DeiT-Tiny-Distil.: We plot the Top-1 Accuracy in the ImageNet-1k validation set for each of the pruning methods as a function of computational efficiency, in the top row measured by GFLOPs and in the bottom row measured by throughput, for multi-layer pruning. For weaker encoders and/or smaller transformer blocks such as DeiT-Tiny, we rather use MLP, which means less parallelism. The throughput of models in multi-layer pruning reaches approximately 5000 images per second, potentially limited by hardware bottlenecks.
Method Acc. GFLOPs TP (imgs/s)
Deit-T-Distil. 74.05 1.27 2607
K = 0.8 Random Drop 72.36 1.06 -
Top-K [15] 73.05 1.04 3023
Zero-TP [21] 73.25 1.06 2319
DynamicViT [17] 71.59 1.08 2805
ToMe [5] 72.99 1.05 2872
TNT+MLP (ours) 73.46 1.07 2938
K = 0.7 Random Drop 71.17 0.96 -
Top-K [15] 71.96 0.94 3295
Zero-TP [21] 72.03 0.97 2438
DynamicViT [17] 71.15 0.98 3031
ToMe [5] 71.06 0.94 3120
TNT+MLP (ours) 72.77 0.98 2939
K = 0.6 Random Drop 69.23 0.86 -
Top-K [15] 70.28 0.84 3815
Zero-TP [21] 69.99 0.87 2677
DynamicViT [17] 70.36 0.88 3476
ToMe [5] 66.56 0.84 3624
TNT+MLP (ours) 71.70 0.88 3142
K = 0.5 Random Drop 66.50 0.77 -
Top-K [15] 67.74 0.74 4270
Zero-TP [21] 67.31 0.78 2630
DynamicViT [17] 68.82 0.79 3851
ToMe [5] 57.52 0.75 4043
TNT+MLP (ours) 70.09 0.79 3210
K = 0.45 Random Drop 64.43 0.73 -
Top-K [15] 65.78 0.69 4484
Zero-TP [21] 65.54 0.73 2519
DynamicViT [17] 67.69 0.74 4016
TNT+MLP (ours) 68.43 0.74 3315
Table 13: Single-layer pruning for DeiT-T-Distil. in low-token regime.
Method Acc. GFLOPs TP (imgs/s)
Deit-T-Distil. 74.05 1.27 2607
K = 0.4 Random Drop 61.88 0.68 -
Top-K [15] 63.27 0.65 4814
Zero-TP [21] 63.29 0.69 2864
DynamicViT [17] 66.26 0.70 4293
TNT+MLP (ours) 66.52 0.70 3267
K = 0.35 Random Drop 57.99 0.64 -
Top-K [15] 60.13 0.60 5103
Zero-TP [21] 60.21 0.64 2728
DynamicViT [17] 63.85 0.65 4510
TNT+MLP (ours) 63.89 0.65 3346
K = 0.3 Random Drop 52.65 0.59 -
Top-K [15] 55.79 0.55 5242
Zero-TP [21] 56.20 0.60 2673
DynamicViT [17] 61.03 0.60 4676
TNT+MLP (ours) 60.08 0.61 3512
K = 0.25 Random Drop 46.18 0.55 -
Top-K [15] 50.67 0.51 5108
Zero-TP [21] 51.39 0.56 2598
DynamicViT [17] 57.03 0.56 5016
TNT+MLP (ours) 54.98 0.57 3477
K = 0.2 Random Drop 37.29 0.51 -
Top-K [15] 43.40 0.46 5153
Zero-TP [21] 44.04 0.51 2694
DynamicViT [17] 50.62 0.52 5152
TNT+MLP (ours) 46.77 0.52 3725
Table 14: Single-layer pruning for DeiT-T-Distil. in low-token regime.
Method Acc. GFLOPs TP (imgs/s) Param.
Deit-T-Distil. 74.05 1.27 2607 -
GFLOPs $\approx 0.93$ Top-K [15] 72.46 0.93 3416 Rate=[.9, .9, .8]
Zero-TP [21] 72.45 0.93 2048 Rate=[1., .9, .9, .9, 1.]
DynamicViT [17] 71.61 0.94 3246 $\rho$=.8
ToMe [5] 73.52 0.94 2424 $r$=8
TNT+MLP (ours) 72.82 0.96 2947 Rate=[1., .95, .95]
GFLOPs $\approx 0.82$ Top-K [15] 70.88 0.83 3799 Rate=[.85, .8, .8]
Zero-TP [21] 70.36 0.80 2219 Rate=[1., .8, .8, .9, 1.]
DynamicViT [17] 70.82 0.80 3668 $\rho$=.7
ToMe [5] 72.91 0.82 2401 $r$=11
TNT+MLP (ours) 71.83 0.83 3470 Rate=[.9, .9, .9]
GFLOPs $\approx 0.7$ Top-K [15] 68.18 0.72 4269 Rate=[.85, .7, .7]
Zero-TP [21] 66.24 0.69 2264 Rate=[1., .7, .7, .8, 1.]
DynamicViT [17] 68.63 0.69 4220 $\rho$=.6
ToMe [5] 71.19 0.67 2368 $r$=15
TNT+MLP (ours) 70.36 0.73 3879 Rate=[.85, .85, .8]
GFLOPs $\approx 0.62$ Top-K [15] 63.80 0.62 4751 Rate=[.7, .7, .65]
Zero-TP [21] 60.48 0.61 2367 Rate=[1., .6, .7, .7, 1.]
DynamicViT [17] 64.55 0.60 4715 $\rho$=.5
TNT+MLP (ours) 67.45 0.64 4332 Rate=[.75, .85, .7]
GFLOPs $\approx 0.58$ Top-K [15] 61.17 0.59 5092 Rate=[.65, .65, .65]
Zero-TP [21] 56.84 0.59 2380 Rate=[1., .6, .6, .7, 1.]
DynamicViT [17] 60.34 0.56 4953 $\rho$=.45
TNT+MLP (ours) 64.48 0.59 4304 Rate=[.65, .85, .7]
Table 15: Multi-layer pruning for DeiT-T-Distil. in high-token regime.
Method Acc. GFLOPs TP (imgs/s) Param.
Deit-T-Distil. 74.05 1.27 2607 -
GFLOPs $\approx 0.53$ Top-K [15] 56.15 0.53 4933 Rate=[.6, .6, .6]
Zero-TP [21] 51.41 0.56 2280 Rate=[1., .5, .5, .6, 1.]
DynamicViT [17] 55.20 0.53 4790 $\rho$=.4
ToMe [5] 61.20 0.52 2386 $r$=20
TNT+MLP (ours) 61.59 0.55 4294 Rate=[.6, .8, .7]
GFLOPs $\approx 0.5$ Top-K [15] 51.39 0.49 4987 Rate=[.6, .5, .6]
Zero-TP [21] 35.26 0.51 2355 Rate=[1., .4, .4, .4, 1.]
DynamicViT [17] 47.87 0.50 4448 $\rho$=.35
ToMe [5] 53.77 0.48 2382 $r$=21
TNT+MLP (ours) 56.85 0.51 4314 Rate=[.5, .8, .7]
GFLOPs $\approx 0.47$ Top-K [15] 48.70 0.47 5064 Rate=[.55, .5, .6]
Zero-TP [21] 24.42 0.49 2298 Rate=[1., .4, .3, .4, 1.]
DynamicViT [17] 36.48 0.47 4823 $\rho$=.3
ToMe [5] 51.15 0.46 2366 $r$=22
TNT+MLP (ours) 53.58 0.48 4368 Rate=[.5, .7, .7]
GFLOPs $\approx 0.45$ Top-K [15] 44.47 0.46 5031 Rate=[.55, .5, .5]
Zero-TP [21] 22.97 0.46 2339 Rate=[1., .3, .4, .4, 1.]
DynamicViT [17] 24.86 0.44 4823 $\rho$=.25
ToMe [5] 44.34 0.44 2369 $r$=24
TNT+MLP (ours) 50.60 0.46 4421 Rate=[.5, .7, .6]
GFLOPs $\approx 0.43$ Top-K [15] 39.54 0.43 4919 Rate=[.5, .45, .5]
Zero-TP [21] 6.97 0.45 2374 Rate=[1., .3, .3, .4, 1.]
DynamicViT [17] 9.65 0.42 4292 $\rho$=.2
ToMe [5] 39.05 0.43 2382 $r$=25
TNT+MLP (ours) 44.41 0.44 4374 Rate=[.5, .6, .55]
Table 16: Multi-layer pruning for DeiT-T-Distil. in low-token regime.
B.4 Base Model: ViT/16
For ViT/16, we use the mean-pooled tokens for the prediction rather than using CLS token. Therefore, Top-K is not applicable. We instead sweep the number of tokens (denoted as #tokens) from 160 to 50. #Tokens from 160 to 110 represent the high-token regime while #tokens from 100 to 50 correspond to low-token regime. For single-layer pruning, high-token regime results are listed in Table 17; low-token regime results are listed in Table 18. For multi-layer pruning, high-token regime results are listed in Table 19; low-token regime results are listed in Table 20.
#Tokens Method Acc. GFLOPs TP (imgs/s)
ViT/16 78.70 9.17 644
160 Random Drop 77.17 7.69 -
Zero-TP [21] 77.43 7.32 665
ToMe [5] 77.71 7.66 752
TNT (ours) 77.94 7.70 747
150 Random Drop 76.75 7.29 -
Zero-TP [21] 76.80 6.92 687
ToMe [5] 77.22 7.24 784
TNT (ours) 77.64 7.30 779
140 Random Drop 76.00 6.89 -
Zero-TP [21] 76.80 6.92 722
ToMe [5] 76.53 6.83 838
TNT (ours) 76.97 6.90 820
130 Random Drop 75.12 6.50 -
Zero-TP [21] 75.96 6.52 749
ToMe [5] 75.70 6.42 877
TNT (ours) 76.22 6.50 856
120 Random Drop 73.79 6.10 -
Zero-TP [21] 74.81 6.13 812
ToMe [5] 74.19 6.02 969
TNT (ours) 75.24 6.11 945
110 Random Drop 71.95 5.71 -
Zero-TP [21] 73.26 5.74 863
ToMe [5] 71.54 5.62 1047
TNT (ours) 73.73 5.72 1016
Table 17: Single-layer pruning for ViT/16. in high-token regime.
#Tokens Method Acc. GFLOPs TP (imgs/s)
ViT/16 78.70 9.17 644
100 Random Drop 69.21 5.33 -
Zero-TP [21] 70.92 5.35 901
ToMe [5] 66.46 5.22 1103
TNT (ours) 71.69 5.33 1069
90 Random Drop 64.71 4.94 -
Zero-TP [21] 67.40 4.97 960
TNT (ours) 68.73 4.95 1153
80 Random Drop 57.64 4.56 -
Zero-TP [21] 61.76 4.59 1002
TNT (ours) 63.87 4.56 1247
70 Random Drop 46.28 4.18 -
Zero-TP [21] 53.15 4.21 1021
TNT (ours) 55.19 4.19 1344
60 Random Drop 31.16 3.81 -
Zero-TP [21] 40.84 3.83 1157
TNT (ours) 41.75 3.81 1457
50 Random Drop 16.57 3.44 -
Zero-TP [21] 27.64 3.46 1228
TNT (ours) 27.40 3.44 1531
Table 18: Single-layer pruning for ViT/16. in low-token regime.
Method Acc. GFLOPs TP (imgs/s) Param.
ViT/16 78.70 9.17 644 -
GFLOPs $\approx 6.8$ Zero-TP [21] 77.07 6.75 688 Rate=[1., .9, .9, .9, 1.]
ToMe [5] 77.02 6.97 734 $r$=8
TNT (ours) 77.32 6.73 842 Rate=[1., .95, .95]
GFLOPs $\approx 6.00$ Zero-TP [21] 75.16 5.84 778 Rate=[1., .8, .8, .9, 1.]
ToMe [5] 77.02 6.14 823 $r$=11
TNT (ours) 75.47 5.84 967 Rate=[.9, .9, .9]
GFLOPs $\approx 5.00$ Zero-TP [21] 69.91 4.98 885 Rate=[1., .7, .7, .8, 1.]
ToMe [5] 70.95 5.06 989 $r$=15
TNT (ours) 72.23 5.10 1103 Rate=[.85, .85, .8]
GFLOPs $\approx 4.43$ Zero-TP [21] 60.39 4.41 966 Rate=[1., .6, .7, .7, 1.]
TNT (ours) 65.81 4.46 1249 Rate=[.75, .85, .7]
GFLOPs $\approx 4.10$ Zero-TP [21] 51.96 4.25 994 Rate=[1., .6, .6, .7, 1.]
TNT (ours) 58.96 4.08 1351 Rate=[.65, .85, .7]
Table 19: Multi-layer pruning for ViT/16. in high-token regime.
Method Acc. GFLOPs TP (imgs/s) Param.
ViT/16 78.70 9.17 644 -
GFLOPs $\approx 3.90$ Zero-TP [21] 40.38 3.98 1035 Rate=[1., .6, .6, .7, 1.]
ToMe [5] 41.78 3.96 1243 $r$=20
TNT (ours) 51.40 3.80 1452 Rate=[.65, .85, .7]
GFLOPs $\approx 3.60$ Zero-TP [21] 13.86 3.60 1092 Rate=[1., .5, .5, .6, 1.]
ToMe [5] 22.58 3.64 1340 $r$=21
TNT (ours) 39.81 3.50 1546 Rate=[.5, .8, .7]
GFLOPs $\approx 3.40$ Zero-TP [21] 5.15 3.48 1102 Rate=[1., .4, .3, .4, 1.]
ToMe [5] 17.74 3.51 1382 $r$=21
TNT (ours) 30.86 3.31 1626 Rate=[.5, .7, .7]
GFLOPs $\approx 3.20$ Zero-TP [21] 4.12 3.27 1163 Rate=[1., .3, .4, .4, 1.]
ToMe [5] 9.19 3.37 1446 $r$=24
TNT (ours) 23.71 3.18 1700 Rate=[.5, .7, .6]
GFLOPs $\approx 3.10$ Zero-TP [21] 0.64 3.17 1179 Rate=[1., .3, .3, .4, 1.]
ToMe [5] 6.30 3.26 1485 $r$=25
TNT (ours) 13.42 2.98 1786 Rate=[.5, .6, .55]
Table 20: Multi-layer pruning for ViT/16. in low-token regime.
C More examples for visualization
More qualitative results for pruning $50\%$ of tokens at varying layers (single-layer pruning at layers 1-5) on the ImageNet-1K validation dataset in Figure  8.
Figure 8: More Visualization of Token Pruning maps on ImageNet-1K: at left are the original images, and at each column progressing right are single layer prunings and their associated kept/dropped tokens, for layers 1-5 of the DeiT-S-Distil. model.